1. About what percent of data is within 1 standard deviation of the mean?


2. About what percent of data is within the mean and 2 standard deviations above the mean?


3. About what percent of data is more than 2 standard deviations above the mean?
1
Expert's answer
2020-04-16T16:29:07-0400

when consider a normal distribution with mean =0 and standard deviation=1 and z is the z value in standard notation.

1) percent of data is within 1 standard deviation of the mean"(P_1),"

"P_1=P(-1<z<1)=2*P(0<z<1)"

(using calculator or table probability can be calculated)

"P_1=2*0.34134=0.68268=\bold{68.268\%}"

2) percent of data is within the mean and 2 standard deviations above the mean"(P_2)" ,

"P_2=P(0<z<2) \\"

(using calculator or table probability can be calculated)

"P_2=0.47725=\bold{47.725\%}"

3) percent of data is more than 2 standard deviations above the mean "(P_3),"

"P_3=P(z>2)\\"

(using calculator or table probability can be calculated)

"P_3=0.02275=\bold{2.275\%}"